OpenStudy (anonymous):

how do u simplify 3x+2y=8} -2y=x

OpenStudy (anonymous):

This is nonsense...

OpenStudy (nilankshi):

ya

OpenStudy (nilankshi):

rite

OpenStudy (anonymous):

cannot be simplified at the simplest

OpenStudy (radar):

3x+2y=8 3(-2y)+2y=8 -6y+2y=8 -4y=8 y=-2 then x would be, x=-2y=-2(-2)=4

OpenStudy (radar):

A solution, not a simplification.